Suppose That A Flare Is Launched Upward With An Initial Velocity Of 80 Ft/sec From A Height Of 224 Ft. Its Height In Feet, H(t), After T Seconds Is Given By H(t) = -16t2 + 80t + 224 After How Long Will The Flare Reach The Ground?

Answer by ankor@dixie-net.com(15622)

Suppose That A Flare Is Launched Upward With An Initial Velocity Of 80 Ft/sec From A Height Of 224 Ft. Its Height In Feet, H(t), After T Seconds Is Given By H(t) = -16t2 + 80t + 224 After How Long Will The Flare Reach The Ground?
:
When the flare reaches the ground, H = 0, find t:
:
-16t^2 + 80t + 224 = 0
we can greatly simplify this, divide by -16
t^2 - 5t - 14 = 0
This factors to
(t-7)(t+2) = 0
Positive solution
t = 7 seconds after it is launched
;
;
Check this; replace x with 7 in the original equation
H = -16(7^2) + 80(7) + 224
H = -16(49) + 560 + 224
H = -784 + 784
H = 0